# The Central Role of the Loss Function in Reinforcement Learning

### v2

@@ Abstract
First-Order and Second-Order Bounds,
@@ 1 Introduction
Table 1: The rate of convergence in decision-making regret achievable by each loss function, where $n$ is the number of samples or interactions. We see that the squared loss is not able to adapt to small-cost or small-variance settings while the mle loss can.
@@ 2.1 Solving CSC with Squared-Loss Regression
A value-based approach to CSC is to learn a cost prediction $f(x,a)\approx\bar{C}(x,a)$ by regressing costs on contexts and then use an induced policy: $\pi_{f}(x)\in\mathop{\rm arg\,min}_{a}f(x,a)$. A standard way to learn such a cost prediction is to minimize squared error [5, 43, 28].
Define the squared loss as
$\ell_{\operatorname{sq}}(\hat{y},y):=(\hat{y}-y)^{2},$
and define the excess squared-loss risk of a prediction $f$ as $\mathcal{E}_{\operatorname{sq}}(f):=\sum_{a}\mathbb{E}[\ell_{\operatorname{sq}}(f(x,a),c(a))-\ell_{\operatorname{sq}}(\bar{C}(x,a),c(a))]$. This straightforwardly bounds the suboptimality of its induced policy:
@@
Under Assump. 1, for any $\delta\in(0,1)$, with probability at least (w.p.a.l.) $1-\delta$,
$\mathcal{E}_{\operatorname{sq}}(\hat{f}^{\operatorname{sq}}_{\mathcal{F}})\lesssim A\log(\left\lvert\mathcal{F}\right\rvert/\delta)/n.$
@@
Under Assump. 1, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in squared loss regression enjoys
$V(\pi_{\hat{f}^{\operatorname{sq}}_{\mathcal{F}}})-V^{\star}\lesssim\sqrt{A\log(\left\lvert\mathcal{F}\right\rvert/\delta)/n}.$
@@ 2.2 The Second-Order Lemma
Here, $h^{2}(p,q)$ is the squared Hellinger distance, which is an $f$-divergence, and it is bounded in $[0,1]$. This lemma is equivalent to Lemma 4.3 of [62] and we provide a simplified proof in Sec. 2.6. Interpreting the inequality, which is a completely deterministic statement, in terms of estimating means, it says that estimation error can be bounded by two terms: one involves the standard deviation times a discrepancy and the other is a squared discrepancy. As variance shrinks, the first term vanishes and the second term dominates, which as a squared term we expect to decay quickly. We note [24, 25] are related and provide first-moment and second-moment bounds, which do not directly imply the tighter variance bound in Lem. 1.
@@ 2.3 Regression with the Binary-Cross-Entropy Loss: First-Order PAC Bounds for CSC
For the final part of the proof, we need to show that minimizing empirical bce-loss risk gives good control on $\mathcal{E}_{\operatorname{bce}}(\hat{f}^{\operatorname{bce}}_{\mathcal{F}})$. We do so with the following symmetrization lemma.
@@
Applying the lemma with $Z_{i}=\frac{1}{2}\ell_{\operatorname{bce}}(f(x_{i},a),c_{i}(a))-\frac{1}{2}\ell_{\operatorname{bce}}(\bar{C}(x_{i},a),c_{i}(a))$ with union bound over $a\in\mathcal{A}$ and $f\in\mathcal{F}$, we have w.p.a.l. $1-\delta$, for all $f\in\mathcal{F}$
@@
$\textstyle-\ell_{\operatorname{bce}}(\bar{C}(x_{i},a),c_{i}(a))+A\ln(2\left\lvert\mathcal{F}\right\rvert/\delta).$
With Assump. 1, the empirical minimizer $\hat{f}^{\operatorname{bce}}_{\mathcal{F}}$ enjoys
@@
Under Assump. 1, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in bce loss regression enjoys
@@
A refinement of Eq. 2 keeps the first term as $\sqrt{f(1-f)}h_{\operatorname{Ber}}$ instead of $\sqrt{f}h_{\operatorname{Ber}}$. This would imply a more refined first-order bound that scales as $\widetilde{\mathcal{O}}(\sqrt{V^{\star}(1-V^{\star})\cdot\frac{1}{n}}+\frac{1}{n})$, where the leading term vanishes also if $V^{\star}\approx 1$. Bounds scaling with $1-V^{\star}$ are sometimes called ‘small-reward’ bounds [6] and are easier to obtain than ‘small-cost’ bounds [44, 61], which we focus on in this paper.
@@ 2.4 Maximum Likelihood Estimation: Second-Order PAC Bounds for CSC
Assumption 2 (Distribution Realizability).
@@
Finally, we apply the symmetrization lemma (Lem. 3) with $Z_{i}=\frac{1}{2}\ell_{\operatorname{mle}}(p(x_{i},a),c_{i}(a))-\frac{1}{2}\ell_{\operatorname{mle}}(C(x_{i},a),c_{i}(a))$ with union bound over $\mathcal{P}$, to deduce that w.p.a.l. $1-\delta$, for all $p\in\mathcal{P}$:
@@
Together with Assump. 2, we have that
@@
Under Assump. 2, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in mle enjoys
@@
Since costs are bounded in $[0,1]$, we observe that $\sigma^{2}(\pi)\leq V(\pi)$, and hence a second-order bound is tighter than a first-order bound. However, our approach is distributional with bounds depending on $\ln|\mathcal{P}|$ which can be larger than $\ln|\mathcal{F}|$; also, Assump. 2 may be more stringent than Assump. 1, although in practice, distributional approaches still achieve superior performance [61, 62, 10].
@@ 2.5 Improved Second-Order PAC Bounds for CSC with Pessimistic MLE
Under Assump. 2, for any $\delta\in(0,1)$, set $\beta=2A\ln(A\left\lvert\mathcal{P}\right\rvert/\delta)$ and define,
@@
where the first inequality is by Eq. 5 and the fact that $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$ minimizes the empirical risk; and the second inequality is by the definition of $\mathcal{P}_{n}$. To prove Claim (b), we first show that $C\in\mathcal{P}_{n}$: by Eq. 5 and the non-negativity of $\mathcal{E}_{\operatorname{mle}}$, we have $L_{\operatorname{mle}}(C)-L_{\operatorname{mle}}(p)\leq 2A\ln(A|\mathcal{P}|/\delta)=\beta$ for all $p\in\mathcal{P}$ (which includes $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$). Thus, this shows that $C$ satisfies the $\mathcal{P}_{n}$ condition, implying its membership in the set. To conclude Claim (b), we have $\sum_{i=1}^{n}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},\pi^{\star}(x_{i}))\geq\sum_{i=1}^{n}\min_{a}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},a)\geq\sum_{i=1}^{n}\min_{a}\bar{C}(x_{i},a)$. Claim (b) then follows by multiplicative Chernoff [71, Theorem 13.5]. ∎
@@
Under Assump. 2, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, pessimistic mle enjoys
@@
$\textstyle\lesssim\sqrt{\sigma^{2}(\widehat{\pi}^{\operatorname{op}})\cdot\frac{A\ln(A|\mathcal{P}|/\delta)}{n}}+\frac{A\ln(A|\mathcal{P}|/\delta)}{n}.$
@@ 2.6 Proof of the Second-Order Lemma
which is a simple consequence of Cauchy-Schwartz [61, Lemma A.1]. Thus, we first prove the second-order lemma using ($\triangle$($\cdot$)△ ( ⋅ ), which is more natural, and then convert the bounds to $h^{2}(\cdot)$ using Eq. 9.
@@
$\overset{ii}{\leq}\textstyle\sqrt{\int_{y}(y-\bar{q})^{2}(p(y)+q(y))\mathrm{d}\lambda^{\prime}(y)\cdot\triangle(p,q)}$
$\overset{iii}{\leq}\textstyle\frac{1}{2}\left(\int_{y}(y-\bar{q})^{2}p(y)\mathrm{d}\lambda^{\prime}(y)+\sigma^{2}(g)\right)+\frac{\triangle(p,q)}{2},$
where (i) is by Cauchy-Schwarz, (ii) is by the premise that $p,q$ are densities on $[0,1]$, and (iii) is by AM-GM. Rearranging terms, we get
$\int_{y}(y-\bar{q})^{2}p(y)\mathrm{d}\lambda^{\prime}(y)\leq 3\sigma^{2}(g)+\triangle(p,q).$
Finally, plugging back into (ii) implies Eq. 10.
@@ 3.1 Plug-in Squared Loss Cannot Achieve First-Order
Probability of the bad event. For $\mathfrak{E}_{1}$, since $n(x^{2})\sim\operatorname{Bin}(n,n^{-1})$, thus $\Pr(n(x^{2})=1)=(1-n^{-1})^{n-1}\geq e^{-1}$. Hence, $\Pr(\mathfrak{E}_{1})\geq(2e)^{-1}$. For $\mathfrak{E}_{2}$, we apply the multiplicative Chernoff bound [71, Theorem 13.5], which implies $\widehat{\mu}(x^{1},a^{1})<2\mu_{n}+\frac{3}{n-1}$ w.p.a.l. $1-e^{-3}$. Thus, $\Pr(\mathfrak{E}_{1}\cap\mathfrak{E}_{2})\geq 1-(1-(2e)^{-1})-e^{-3}\geq 0.1$. ∎
@@ 3.2 Plug-in BCE Loss Cannot Achieve Second-Order
4 Reinforcement Learning

### v3

@@ Abstract
First-Order (Small-Loss) and Second-Order (Variance-Dependent) Bounds,
@@ 1 Introduction
Table 1: The decision-making regret of value-based RL per loss function, where $n$ is the number of samples. In this paper, we’ll see that squared loss cannot adapt to small-cost or small-variance settings while binary-cross-entropy (bce) and maximum likelihood estimation (mle) can. We remark that mle is used with distributional regression and thus requires stronger realizability or completeness conditions. Moreover, all three losses use slightly different eluder dimensions in the online setting. We summarize these nuances in Table 2.
@@ 2.1 Solving CSC with Squared-Loss Regression
A value-based approach to CSC is to learn a cost prediction $f(x,a)\approx\bar{C}(x,a)$ by regressing costs on contexts and then use an induced greedy policy: $\pi_{f}(x)\in\mathop{\rm arg\,min}_{a}f(x,a)$. A standard way to learn such a cost prediction is to minimize the squared error [5, 45, 30].
Define the squared loss and the excess squared-loss risk for a prediction function $f$ as:
$\textstyle\ell_{\operatorname{sq}}(\hat{y},y):=(\hat{y}-y)^{2},$
$\textstyle\mathcal{E}_{\operatorname{sq}}(f):=\sum_{a}\mathbb{E}[\ell_{\operatorname{sq}}(f(x,a),c(a))-\ell_{\operatorname{sq}}(\bar{C}(x,a),c(a))].$
This can be used to expediently bound the sub-optimality of the policy induced by $f$:
@@
Under Asm. 1, for any $\delta\in(0,1)$, with probability at least (w.p.a.l.) $1-\delta$,
$\mathcal{E}_{\operatorname{sq}}(\hat{f}^{\operatorname{sq}}_{\mathcal{F}})\lesssim A\log({A}\left\lvert\mathcal{F}\right\rvert/\delta)/n.$
@@
Under Asm. 1, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in squared loss regression enjoys
$V(\pi_{\hat{f}^{\operatorname{sq}}_{\mathcal{F}}})-V^{\star}\lesssim\sqrt{A\log({A}\left\lvert\mathcal{F}\right\rvert/\delta)/n}.$
@@ 2.2 The Second-Order Lemma
Here, $h^{2}(p,q)$ is the squared Hellinger distance, which is an $f$-divergence, and it is bounded in $[0,1]$. This lemma is equivalent to Lemma 4.3 of [65] and we provide a simplified proof in Sec. 2.6. Interpreting the inequality, which is a completely deterministic statement, in terms of estimating means, it says that estimation error can be bounded by two terms: one involves the standard deviation times a discrepancy and the other is a squared discrepancy. As variance shrinks, the first term vanishes and the second term dominates, which as a squared term we expect to decay quickly. We note that [26, 27] obtain related first-moment and second-moment bounds, which are generally looser and do not directly imply the variance bound in Lem. 1.
@@ 2.3 Regression with the Binary-Cross-Entropy Loss: First-Order PAC Bounds for CSC
For the final part of the proof, we need to show that minimizing the empirical bce-loss risk gives good control on $\mathcal{E}_{\operatorname{bce}}(\hat{f}^{\operatorname{bce}}_{\mathcal{F}})$. This is implied by the following tail bound.
@@
Then, applying Lem. 3 with $Z_{i}=\frac{1}{2}\ell_{\operatorname{bce}}(f(x_{i},a),c_{i}(a))-\frac{1}{2}\ell_{\operatorname{bce}}(\bar{C}(x_{i},a),c_{i}(a))$ and a union bound over $a$ and $f$, we have w.p.a.l. $1-\delta$, for all $f\in\mathcal{F}$
@@
$\textstyle-\ell_{\operatorname{bce}}(\bar{C}(x_{i},a),c_{i}(a))+A\ln(2A\left\lvert\mathcal{F}\right\rvert/\delta).$
With Asm. 1, the empirical minimizer $\hat{f}^{\operatorname{bce}}_{\mathcal{F}}$ enjoys
@@
Under Asm. 1, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in bce loss regression enjoys
@@
A refinement of Eq. 2 keeps the first term as $\sqrt{f(1-f)}h_{\operatorname{Ber}}$ instead of $\sqrt{f}h_{\operatorname{Ber}}$. This would imply a more refined first-order bound that scales as $\widetilde{\mathcal{O}}(\sqrt{V^{\star}(1-V^{\star})\cdot\frac{1}{n}}+\frac{1}{n})$, where the leading term vanishes also if $V^{\star}\approx 1$. Bounds scaling with $1-V^{\star}$ are sometimes called ‘small-reward’ bounds [6], which are generally easier to prove than ‘small-cost’ bounds [46, 64]. For example, if the bound scales with the minimum achievable reward $R^{\star}$, the case where $R^{\star}=0$ is actually trivial since all policies are optimal. However, if the bound scales with the minimum achievable cost $L^{\star}$, the case where $L^{\star}=0$ is interesting because sub-optimal policies may still have large cost. Thus, following Lykouris, Sridharan and Tardos [46], Wang et al. [64], we focus on the cost minimizing setting in this paper.
@@ 2.4 Maximum Likelihood Estimation: Second-Order PAC Bounds for CSC
Assumption 2 (Distributional Realizability).
@@
While this assumption is more stringent than mean-realizability (Asm. 1 from before), it is required to prove that MLE succeeds [20, 3]. Alternative algorithms such as Scheffé tournament can be used to learn distributions under misspecification, but they are computationally hard to implement [2].
Finally, we apply Lem. 3 with $Z_{i}=\frac{1}{2}\ell_{\operatorname{mle}}(p(x_{i},a),c_{i}(a))-\frac{1}{2}\ell_{\operatorname{mle}}(C(x_{i},a),c_{i}(a))$ with union bound over $\mathcal{P}$, to deduce that w.p.a.l. $1-\delta$, for all $p\in\mathcal{P}$:
@@
Together with Asm. 2, we have that
@@
Under Asm. 2, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, plug-in mle enjoys
@@
Since costs are bounded in $[0,1]$, we observe that $\sigma^{2}(\pi)\leq V(\pi)$, and hence a second-order bound is tighter than a first-order bound. However, our approach is distributional with bounds depending on $\ln|\mathcal{P}|$ which can be larger than $\ln|\mathcal{F}|$, and as mentioned earlier, realizability in distribution space is stricter than realizability in the conditional mean. We remark that distributional approaches still achieve superior performance [10, 11, 64, 65] in practice, suggesting that MLE with modern function approximators can well-approximate complex distributions despite the stronger assumption in theory.
@@ 2.5 Improved Second-Order PAC Bounds for CSC with Pessimistic MLE
Under Asm. 2, for any $\delta\in(0,1)$, set $\beta=2A\ln(A\left\lvert\mathcal{P}\right\rvert/\delta)$ and define,
@@
where the first inequality is by Eq. 5 and the fact that $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$ minimizes the empirical risk; and the second inequality is by the definition of $\mathcal{P}_{n}$. To prove Claim (b), we first show that $C\in\mathcal{P}_{n}$: by Eq. 5 and the non-negativity of $\mathcal{E}_{\operatorname{mle}}$, we have $L_{\operatorname{mle}}(C)-L_{\operatorname{mle}}(p)\leq 2A\ln(A|\mathcal{P}|/\delta)=\beta$ for all $p\in\mathcal{P}$ (which includes $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$). Thus, this shows that $C$ satisfies the $\mathcal{P}_{n}$ condition, implying its membership in the set. To conclude Claim (b), we have $\sum_{i=1}^{n}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},\pi^{\star}(x_{i}))\geq\sum_{i=1}^{n}\min_{a}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},a)\geq\sum_{i=1}^{n}\min_{a}\bar{C}(x_{i},a)$. Claim (b) then follows by multiplicative Chernoff [75, Theorem 13.5]. ∎
@@
Under Asm. 2, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, pessimistic mle enjoys
@@
$\textstyle\lesssim\sqrt{{\color[rgb]{1,0,0}\sigma^{2}(\widehat{\pi}^{\operatorname{op}})}\cdot\frac{A\ln(A|\mathcal{P}|/\delta)}{n}}+\frac{A\ln(A|\mathcal{P}|/\delta)}{n}.$
@@ 2.6 Proof of the Second-Order Lemma
which is a simple consequence of Cauchy-Schwartz [64, Lemma A.1]. Thus, we first prove the second-order lemma using ($\triangle$($\cdot$)△ ( ⋅ ), which is more natural, and then convert the bounds to $h^{2}(\cdot)$ using Eq. 9.
@@
$\overset{ii}{\leq}\textstyle\sqrt{\int_{y}(y-\bar{q})^{4}(p(y)+q(y))\mathrm{d}\lambda^{\prime}(y)\cdot\triangle(p,q)}$
$\overset{iii}{\leq}\textstyle\sqrt{\int_{y}(y-\bar{q})^{2}(p(y)+q(y))\mathrm{d}\lambda^{\prime}(y)\cdot\triangle(p,q)}$
$\overset{iv}{\leq}\textstyle\frac{1}{2}\left(\int_{y}(y-\bar{q})^{2}p(y)\mathrm{d}\lambda^{\prime}(y)+\sigma^{2}(q)\right)+\frac{\triangle(p,q)}{2},$
where (i) is by the definition of variance, (ii) is by Cauchy-Schwarz, (iii) is by the premise that $p,q$ are densities on $[0,1]$, and (iv) is by AM-GM. Rearranging the inequality that (i)$\leq$(iv), we get
$\int_{y}(y-\bar{q})^{2}p(y)\mathrm{d}\lambda^{\prime}(y)\leq 3\sigma^{2}(q)+\triangle(p,q).$
Finally, plugging this back into (iii) implies Eq. 10.
@@ 3.1 Plug-in Squared Loss Cannot Achieve First-Order
Probability of the bad event. For $\mathfrak{E}_{1}$, since $n(x^{2})\sim\operatorname{Bin}(n,n^{-1})$, thus $\Pr(n(x^{2})=1)=(1-n^{-1})^{n-1}\geq e^{-1}$. Hence, $\Pr(\mathfrak{E}_{1})\geq(2e)^{-1}$. For $\mathfrak{E}_{2}$, we apply the multiplicative Chernoff bound [75, Theorem 13.5], which implies $\widehat{\mu}(x^{1},a^{1})<2\mu_{n}+\frac{3}{n-1}$ w.p.a.l. $1-e^{-3}$. Thus, $\Pr(\mathfrak{E}_{1}\cap\mathfrak{E}_{2})\geq 1-(1-(2e)^{-1})-e^{-3}\geq 0.1$. ∎
@@ 3.2 Plug-in BCE Loss Cannot Achieve Second-Order
Table 2: Summary of main RL results in this paper. Since the algorithms presented all use temporal difference learning, all results posit some variant of Bellman Completeness. In online RL, each bound uses a slightly different eluder dimension based on the loss; in offline RL, all bounds use the single policy coverage $C^{\widetilde{\pi}}$ (defined in Eq. 21). Finally, we remark that the online and offline RL algorithms employ optimism and pessimism over a version space, so they are computationally inefficient [18]. For computational efficiency, we study the hybrid RL setting (Sec. 8) where we show that fitted-Q iteration with access to online and offline data achieves a bound that is the sum of online and offline bounds.
Setting Loss Function Class Bellman Completeness Regret / PAC Bound
Online RL $\ell_{\operatorname{sq}}$ $\mathcal{F}:\mathcal{X}\times\mathcal{A}\mapsto[0,1]$ $\mathcal{T}^{\star}$ (Asm. 3) $\mathcal{O}\left(H\sqrt{Kd_{\operatorname{sq}}\ln(H|\mathcal{F}|/\delta)}\right)$ (Thm. 7)
$\ell_{\operatorname{bce}}$ $\mathcal{F}:\mathcal{X}\times\mathcal{A}\mapsto[0,1]$ $\mathcal{T}^{\star}$ (Asm. 3) $\mathcal{O}\left(H\sqrt{{\color[rgb]{1,0,0}V^{\star}}\cdot Kd_{\operatorname{bce}}\ln(H|\mathcal{F}|/\delta)}+H^{2}d_{\operatorname{bce}}\ln(H|\mathcal{F}|/\delta)\right)$ (Thm. 8)
$\ell_{\operatorname{mle}}$ $\mathcal{P}:\mathcal{X}\times\mathcal{A}\mapsto\Delta([0,1])$ $\mathcal{T}^{\operatorname{D},\star}$ (Asm. 5) $\mathcal{O}\left(H\sqrt{{\color[rgb]{1,0,0}\sum_{k=1}^{K}\sigma^{2}(\pi^{k})}\cdot d_{\operatorname{mle}}\ln(H|\mathcal{P}|/\delta)}+H^{2.5}d_{\operatorname{mle}}\ln(H|\mathcal{P}|/\delta)\right)$ (Thm. 9)
Offline RL $\ell_{\operatorname{sq}}$ $\mathcal{F}:\mathcal{X}\times\mathcal{A}\mapsto[0,1]$ $\mathcal{T}^{\pi},\forall\pi\in\Pi$ (Asm. 6) $\mathcal{O}\left(H\sqrt{\frac{C^{\widetilde{\pi}}\ln(H|\mathcal{F}|/\delta)}{n}}\right)$ (Thm. 10)
$\ell_{\operatorname{bce}}$ $\mathcal{F}:\mathcal{X}\times\mathcal{A}\mapsto[0,1]$ $\mathcal{T}^{\pi},\forall\pi\in\Pi$ (Asm. 6) $\mathcal{O}\left(H\sqrt{{\color[rgb]{1,0,0}V^{\widetilde{\pi}}}\cdot\frac{C^{\widetilde{\pi}}\ln(H|\mathcal{F}|/\delta)}{n}}+H^{2}\frac{C^{\widetilde{\pi}}\ln(H|\mathcal{F}|/\delta)}{n}\right)$ (Thm. 10)
$\ell_{\operatorname{mle}}$ $\mathcal{P}:\mathcal{X}\times\mathcal{A}\mapsto\Delta([0,1])$ $\mathcal{T}^{\operatorname{D},\pi},\forall\pi\in\Pi$ (Asm. 7) $\mathcal{O}\left(H\sqrt{{\color[rgb]{1,0,0}\sigma^{2}(\widetilde{\pi})}\cdot\frac{C^{\widetilde{\pi}}\ln(H|\mathcal{P}|/\delta)}{n}}+H^{2.5}\frac{C^{\widetilde{\pi}}\ln(H|\mathcal{P}|/\delta)}{n}\right)$ (Thm. 11)
4 Setup for Reinforcement Learning